OpenStudy (anonymous):

Container A contains 990 milliliters of water and 10 milliliters of wine. Container B contains 1995 milliliters of water and 5 milliliters of wine. The contents of both containers are poured into a big bowl and thoroughly mixed. 200 millilters of liquids are then taken out in a cup how much wine in milliliters ends up in the cup? Container A contains 990 milliliters of water and 10 milliliters of wine. Container B contains 1995 milliliters of water and 5 milliliters of wine. The contents of both containers are poured into a big bowl and thoroughly mixed. 200 millilters of liquids are then taken out in a cup how much wine in milliliters ends up in the cup? @Mathematics

OpenStudy (agreene):

990+1995 = 2985 ml water 10+5 = 15 ml wine 2985/15 = 199 water/wine meaning if you take 200 ml solution, 199ml is water 1ml is wine... I'm forgetting everything I know about chemistry and fluid dynamics to answer this incredibly poorly written question >.<